# Schwarz methods with PMLs for Helmholtz problems: fast convergence at high frequency

### v1

@@
Abstract
We discuss parallel (additive) and sequential (multiplicative) variants of overlapping Schwarz methods for the Helmholtz equation in $\mathbb{R}^{d}$, with large real wavenumber and smooth variable wave speed. The radiation condition is approximated by a Cartesian perfectly-matched layer (PML). The domain-decomposition subdomains are overlapping hyperrectangles with Cartesian PMLs at their boundaries. In a recent paper (arXiv:2404.02156), the current authors proved (for both variants) that, after a specified number of iterations – depending on the behaviour of the geometric-optic rays – the error is smooth and smaller than any negative power of the wavenumber $k$. For the parallel method, the specified number of iterations is less than the maximum number of subdomains, counted with their multiplicity, that a geometric-optic ray can intersect. The theory, which is given at the continuous level and makes essential use of semi-classical analysis, assumes that the overlaps of the subdomains and the widths of the PMLs are all independent of the wavenumber. In this paper we extend the results of arXiv:2404.02156 by experimentally studying the behaviour of the methods in the practically important case when both the overlap and the PML width decrease as the wavenumber increases. We find that (at least for constant wavespeed), the methods remain robust to increasing $k$, even for miminal overlap, when the PML is one wavelength wide.
@@ 1 The Helmholtz problem
$-k^{-2}\Delta u-c^{-2}u=f\quad\text{in}\quad\mathbb{R}^{d},$ (1)
with the Sommerfeld radiation condition: $r^{\frac{d-1}{2}}\left(\partial_{r}u-\mathrm{i}ku\right)\rightarrow 0,\ \text{as}\ r=|\mathbf{x}|\rightarrow\infty.$ Here $k\geq 1$ is the wave number, $c\in C^{\infty}(\mathbb{R}^{d})$ is the wavespeed and $f\in L^{2}(\mathbb{R}^{d})$ is the source. While the paper GaGoGrLaSp:24 treats the case of general $d$, in the interests of brevity we restrict here to $d=2$. The method for general $d$ is an obvious generalisation of the one presented here. We assume that both $f$ and $1-c$ are supported in a box $\Omega_{\rm int}:=(0,l)\times(0,d).$
@@
$P_{s}\,u\ :=\ (-k^{-2}\Delta_{s}-c^{-2})\,u$ $=f\quad\text{on }\quad\Omega,$ (3)
subject $\text{subject to}\quad u$ $=0\quad\text{on }\quad\partial\Omega.$
@@ 2 The overlapping Schwarz methods with local PMLs
We consider both parallel and sequential overlapping Schwarz methods for (3), using PML as a subdomain boundary condition. Such a strategy was first proposed (without theory) in To:98 , and has received much recent interest both in the overlapping and non-overlapping cases – for reviews see GaGoGrLaSp:24 and GaZh:22 . Recent work on the overlapping case includes LeJu:19 and BoBoDoTo:22 . Our paper GaGoGrLaSp:24 provides the first wavenunber-explicit theory, both for the overlapping case, and for any domain decomposition method for Helmholtz problems with a non-trivial scatterer.
We cover $\Omega_{\rm int}$ by $N$ overlapping sub-rectangles: $\Omega_{{\rm int},j}:=(a^{j},b^{j})\times(c^{j},d^{j}),\ j=1,\ldots,N$, and we let $\delta>0$ denote the minimum overlap parameter, as defined for example in (ToWi:05, , Assumption 3.1). Each $\Omega_{\rm int,j}$ is then extended to a larger subdomain $\Omega_{j}$ by adding a PML of width $\kappa$ along each edge, analogous to the extension of $\Omega_{\rm int}$ to $\Omega$. (The PML width on interior edges could also be different from $\kappa$ see GaGoGrLaSp:24 .)
@@
$g_{1,j}(x_{1}):=\begin{cases}f_{s}(x_{1}-{b}_{j})&\text{if }x_{1}\geq{b}_{j}\\ 0&\text{if }x_{1}\in({a}_{j},{b}_{j})\\ {-}f_{s}({a}_{j}-x_{1})&\text{if }x_{1}\leq{a}_{j}\end{cases},\hskip 8.5359ptg_{2,j}(x_{2}):=\begin{cases}f_{s}(x_{2}-{d}_{j})&\text{if }x_{2}\geq{d}_{j}\\ 0&\text{if }x_{2}\in({c}_{j},{d}_{j})\\ {-}f_{s}({c}_{j}-x_{2})&\text{if }x_{2}\leq{c}_{j}.\end{cases},$
@@
To knit local solutions together, we introduce a partition of unity on $\mathbb{R}^{2}$, denoted $\{\chi_{j}\}_{j=1}^{N}$, such that $\sum_{j}\chi_{j}\equiv 1\ \text{on}\ \mathbb{R}^{2}$, and, in addition,
@@
    subject $\text{subject to}\quad u_{j}^{2n+1}$ $=u_{j,n}^{\rightarrow}\quad\text{on}\quad\partial\Omega_{j},$
@@
    subject $\text{subject to}\quad u_{j}^{2n+2}$ $=u_{j,n}^{\leftarrow}\quad\text{on}\quad\partial\Omega_{j},$
@@ 3 Theoretical results
Suppose the POU $\{\chi_{j}\}$ is $C^{\infty}$ Let $\{u^{n}_{*}\}_{n=0}^{\infty}$ be any sequence of iterates for the additive (=+*=+∗ = +) or multiplicative (=×*$\times$ = ×) Schwarz algorithm above. Then (GaGoGrLaSp:24, , Theorems 1.1-1.4 and 1.6) give conditions that, for any $M>0$ and integer $s\geq 1$, guarantee the existence of ${\mathcal{N}}\in\mathbb{N}$ and $C>0$ (both independent of $f$) such that
@@ 4 Discretisation and numerical results
$a(u,v):=\int_{\Omega}k^{-2}\big{(}(D\nabla u)\cdot\nabla\overline{v}-(\beta\cdot\nabla u)\overline{v}\big{)}-c^{-2}u\overline{v}=\langle f,v\rangle_{L^{2}(\Omega)}.$ (10)
@@
$a_{j}\big{(}\mathfrak{c}_{h,j}^{n},v_{h,j}\big{)}=\big{\langle}f,\mathcal{R}^{T}_{h,j}v_{h,j}\big{\rangle}-a\big{(}u^{n}_{h,j},\mathcal{R}^{T}_{h,j}v_{h,j}\big{)}\quad\text{ for all }v_{h,j}\in V_{h,j}.$
@@
Numerical Experiment.  In this experiment, we explore how the overlap size $\delta$ and the definition of the PML affect the performance of the methods. For discretization, we used finite elements of polynomial order $2$, with $h\sim k^{-1.25}$ (to control the pollution error). In the tests in GaGoGrLaSp:24 (chosen to illustrate the theory), we had set $\delta=\kappa=1/40$, fixed independently of $k$. In this case the number of freedoms in the overlaps and the PMLs increases significantly as $k$ increases, implying a heavy communication load at high-frequency. Here we investigate smaller PML thickness $\kappa$ and smaller overlap $\delta$, thus increasing the practicality of the method. The PML scaling function is $f_{\rm PML}(x)=\frac{a}{3}x^{3}$.
@@
Regarding the POU, for a strip DD we start from functions $\tilde{\chi}_{j}$ supported on $\Omega_{{\rm int},j}$, which take value $1$ in the non-overlapping region and decrease linearly to 00 towards the internal boundary of $\Omega_{{\rm int},j}$. Then the PoU is obtained by normalizing, i.e., $\chi_{j}=\frac{\tilde{\chi}_{j}}{\sum_{j}\tilde{\chi}_{j}}.$ For the checkerboard DD, we used the Cartesian product $\tilde{\chi}_{i}(x)\tilde{\chi}_{j}(y)$ to generate local functions on $\Omega_{{\rm int},j}$ and then normalized them to obtain $\chi_{j}$. The POU used here is not as smooth as is required in the theoretical analysis GaGoGrLaSp:24 . In fact, in the experiments in GaGoGrLaSp:24 we used a smooth PoU and each $\chi_{j}$ was supported in a proper subset of $\Omega_{{\rm int},j}$, bounded away from $\partial\Omega_{{\rm int},j}$.
Table 1 lists the iteration counts for the algorithms to attain 1e-6 reduction of the relative residual. Here “RAS-PML” and “RMS-PML” denote respectively the additive and multiplicative algorithms, each tested on strip and checkerboard DDs. For the RMS-PML and checkerboard DD, we use the exhaustive sweeping method that contains 4 sweeps per iteration (see (GaGoGrLaSp:24, , Figure 1.2)). We observe:
*  The performance is hardly degraded when the overlap is reduced from fixed ($\delta=1/80$) to minimal ($\delta=h$). When $\delta=1/80$, the overlap contains $4,7,10,13,16,19$ layers of elements for $k=100,150,200,250,300,350$. The results show that this communication cost can be removed while hardly affecting the convergence.
*  With $\kappa=\lambda$, the iteration counts are not decreasing as fast with increasing $k$ as when we used fixed PML width. (See the experiments in GaGoGrLaSp:24 illustrating (9).) However convergence rates remain bounded for the range of $k$ tested and in the checkerboard case for RAS-PML, the convergence rate decreases as the wave number increases; see Figure 1.
(a) checkerboard, RAS-PML: $\rho=\frac{\|\mathsf{A(u-u^{N_{x}+N_{y}-1})}\|_{\ell^{2}}}{\|\mathsf{A(u-u^{0})}\|_{\ell^{2}}}$
(b) checkerboard, RMS-PML: $\rho=\frac{\|\mathsf{A(u-u^{4})}\|_{\ell^{2}}}{\|\mathsf{A(u-u^{0})}\|_{\ell^{2}}}$
Figure 1: RAS-PML on Checkerboard: The rate of reduction of the residual against $k$
N 2 4 8
k\$\delta$ $\frac{1}{80}$ 2h h $\frac{1}{80}$ 2h h $\frac{1}{80}$ 2h h
@@
(c) RMS-PML:  Left: strip DD, Right checkerboard DD.
Table 1: Iteration counts (relative residual tol = $10^{-6}$) with $f_{\rm PML}(x):=\frac{a}{3}x^{3}$,  $a=30k$. Note that $\frac{1}{80h}=4,7,10,13,16,19$ for $k=100,150,200,250,300,350$
Acknowledgement JG was supported by EPSRC grant EP/V001760/1. SG was supported by the National Natural Science Foundation of China (Grant number 12201535) and the Guangdong Basic and Applied Basic Research Foundation (Grant number 2023A1515011651). DL was supported by INSMI (CNRS) through a PEPS JCJC grant 2023. EAS was supported by EPSRC grant EP/R005591/1. IGG was supported for several collaborative visits by CUHK Shenzhen.

### v2

@@
11institutetext: Jeffrey Galkowski1, Shihua Gong2, Ivan G. Graham3, David Lafontaine4 and Euan A. Spence3 22institutetext: 1. University College London, 2. The Chinese University of Hong Kong, Shenzhen 3. Department of Mathematical Sciences, University of Bath, 4. Institut de Mathematiques de Toulouse, CNRS.
Schwarz methods with PMLs for Helmholtz problems: fast convergence at high frequency
J. Galkowski    S. Gong    I.G. Graham    D. Lafontaine    E.A. Spence
@@ 1 The Helmholtz problem
$k^{-2}\Delta u-c^{-2}u=f\quad\text{in}\quad\mathbb{R}^{d},$ (1)
with the Sommerfeld radiation condition: $r^{\frac{d-1}{2}}\left(\partial_{r}u-\mathrm{i}ku\right)\rightarrow 0,\ \text{as}\ r=|\mathbf{x}|\rightarrow\infty.$ Here $k\geq 1$ is the wave number, $c\in C^{\infty}(\mathbb{R}^{d})$ is the wavespeed and $f\in L^{2}(\mathbb{R}^{d})$ is the source. While the paper GaGoGrLaSp:24 treats the case of general $d$, in the interests of brevity we restrict here to $d=2$. The method for general $d$ is an obvious generalisation of the one presented here. We assume that both $f$ and 1c1-c are supported in a box $\Omega_{\rm int}:=(0,l)\times(0,d).$
@@
$P_{s}\,u\ =\ (-k^{-2}\Delta_{s}-c^{-2})\,u$ $=f\quad\text{on }\quad\Omega,$ (3)
$\text{subject to}\quad u$ $=0\quad\text{on }\quad\partial\Omega.$
@@ 2 The overlapping Schwarz methods with local PMLs
We consider both parallel and sequential overlapping Schwarz methods for (3), using PML as a subdomain boundary condition. Such a strategy was first proposed (without theory) in To:98 , and has received much recent interest both in the overlapping and non-overlapping cases – for reviews see GaGoGrLaSp:24 and GaZh:22 . Recent work on the overlapping case includes LeJu:19 and BoBoDoTo:22 . Our paper GaGoGrLaSp:24 provides the first wavenumber-explicit theory, both for the overlapping case, and for any domain decomposition method for Helmholtz problems with a non-trivial scatterer.
We cover $\Omega_{\rm int}$ by $N$ overlapping sub-rectangles: $\Omega_{{\rm int},j}:=(a^{j},b^{j})\times(c^{j},d^{j}),\ j=1,\ldots,N$, and we let $\delta>0$ denote the minimum overlap parameter, as defined for example in (ToWi:05, , Assumption 3.1). Each $\Omega_{\rm int,j}$ is then extended to a larger subdomain $\Omega_{j}$ by adding a PML of width $\kappa$ along each edge, analogous to the extension of $\Omega_{\rm int}$ to $\Omega$. (The PML width on interior edges could also be different from $\kappa$ - see GaGoGrLaSp:24 .)
@@
$g_{1,j}(x_{1}):=\begin{cases}f_{s}(x_{1}-{b}_{j})&\text{if }x_{1}\geq{b}_{j}\\ 0&\text{if }x_{1}\in({a}_{j},{b}_{j})\\ {-}f_{s}({a}_{j}-x_{1})&\text{if }x_{1}\leq{a}_{j}\end{cases},\hskip 8.5359ptg_{2,j}(x_{2}):=\begin{cases}f_{s}(x_{2}-{d}_{j})&\text{if }x_{2}\geq{d}_{j}\\ 0&\text{if }x_{2}\in({c}_{j},{d}_{j})\\ {-}f_{s}({c}_{j}-x_{2})&\text{if }x_{2}\leq{c}_{j}.\end{cases}.$
@@
To knit local solutions together, we introduce a partition of unity (POU) on $\mathbb{R}^{2}$, denoted $\{\chi_{j}\}_{j=1}^{N}$, such that $\sum_{j}\chi_{j}\equiv 1\ \text{on}\ \mathbb{R}^{2}$, and, in addition,
@@
    $\text{subject to}\quad u_{j}^{2n+1}$ $=u_{j,n}^{\rightarrow}\quad\text{on}\quad\partial\Omega_{j},$
@@
    $\text{subject to}\quad u_{j}^{2n+2}$ $=u_{j,n}^{\leftarrow}\quad\text{on}\quad\partial\Omega_{j},$
@@ 3 Theoretical results
Suppose the POU $\{\chi_{j}\}$ is $C^{\infty}$. Let $\{u^{n}_{*}\}_{n=0}^{\infty}$ be any sequence of iterates for the additive (=+*=+) or multiplicative ($\times$) Schwarz algorithm above. Then (GaGoGrLaSp:24, , Theorems 1.1-1.4 and 1.6) give conditions that, for any $M>0$ and integer $s\geq 1$, guarantee the existence of $\mathcal{N}$$\in\mathbb{N}$ and $C>0$ (both independent of $f$) such that
@@ 4 Discretisation and numerical results
$a(u,v):=\int_{\Omega}k^{-2}\big((D\nabla u)\cdot\nabla\overline{v}-(\beta\cdot\nabla u)\overline{v}\big)-c^{-2}u\overline{v}=\langle f,v\rangle_{L^{2}(\Omega)},$ (10)
@@
$a_{j}\big(\mathfrak{c}_{h,j}^{n},v_{h,j}\big)=\big\langle f,\mathcal{R}^{T}_{h,j}v_{h,j}\big\rangle-a\big(u^{n}_{h,j},\mathcal{R}^{T}_{h,j}v_{h,j}\big)\quad\text{ for all }v_{h,j}\in V_{h,j},$
@@
Numerical Experiment. In this experiment, we explore how the overlap size $\delta$ and the definition of the PML affect the performance of the methods. For simplicity, we consider constant wavespeed $c=1$. For discretization, we used finite elements of polynomial order $2$, with $h\sim k^{-1.25}$ (to control the pollution error). In the tests in GaGoGrLaSp:24 (chosen to illustrate the theory), we had set $\delta=\kappa=1/40$, fixed independently of $k$. In this case the number of freedoms in the overlaps and the PMLs increases significantly as $k$ increases, implying a heavy communication load at high-frequency. Here we investigate smaller PML thickness $\kappa$ and smaller overlap $\delta$, thus increasing the practicality of the method. The PML scaling function is $f_{\rm PML}(x)=\frac{a}{3}x^{3}$.
We tested two possible choices of $\kappa$ and $a$. In the first setting we chose $\kappa=\lambda:=\frac{2\pi}{k}$ (i.e., the PML is one wavelength wide) and we compared different choices of $a\in\{5\times 10^{3},\,10^{4},\,30k,\,0.3k^{2}\}$. In the second setting we chose $\kappa\in\{3h,5h\}$ and $a=k^{2.5}$, (thus ensuring that the PML scaling functions have the same maximal modulus $\mathcal{O}(1)$ in the PML regions). The second setting turned out not to be $k$-robust, while the first was found to perform better when $a$ was increasing with respect to $k$. Due to the page limit, we only present the results for $\kappa=\lambda$ and $a=30k$ and for overlap $\delta\in\{\frac{1}{80},2h,h\}$; when $\delta=h$ the overlap has only two layers of elements.
Regarding the POU, for a strip DD we start from functions $\tilde{\chi}_{j}$ supported on $\Omega_{{\rm int},j}$, which take value $1$ in the non-overlapping region and decrease linearly to 0 towards the internal boundary of $\Omega_{{\rm int},j}$. Then the POU is obtained by normalizing, i.e., $\chi_{j}=\frac{\tilde{\chi}_{j}}{\sum_{j}\tilde{\chi}_{j}}.$ For the checkerboard DD, we used the Cartesian product $\tilde{\chi}_{i}(x)\tilde{\chi}_{j}(y)$ to generate local functions on $\Omega_{{\rm int},j}$ and then normalized them to obtain $\chi_{j}$. The POU used here is not as smooth as is required in the theoretical analysis GaGoGrLaSp:24 . In fact, in the experiments in GaGoGrLaSp:24 we used a smooth POU and each $\chi_{j}$ was supported in a proper subset of $\Omega_{{\rm int},j}$, bounded away from $\partial\Omega_{{\rm int},j}$.
Table 1 lists the iteration counts for the algorithms to attain 1e-6 reduction of the relative residual. Here “RAS-PML” and “RMS-PML” denote respectively the additive and multiplicative algorithms, each tested on strip and checkerboard DDs. For the RMS-PML and checkerboard DD, we use the exhaustive sweeping method that contains 4 sweeps per iteration (see (GaGoGrLaSp:24, , Figure 1.2)). We observe: 1) The performance is hardly degraded when the overlap is reduced from fixed ($\delta=1/80$) to minimal ($\delta=h$). When $\delta=1/80$, the overlap contains $4,7,10,13,16,19$ layers of elements for $k=100,150,200,250,300,350$. The results show that this communication cost can be removed while hardly affecting the convergence. 2) With $\kappa=\lambda$, the iteration counts are not decreasing as fast with increasing $k$ as when we used fixed PML width. (See the experiments in GaGoGrLaSp:24 illustrating (9).) However convergence rates remain bounded for the range of $k$ tested and in the checkerboard case for RAS-PML, the convergence rate decreases as the wave number increases; see Figure 1.
(a) RAS-PML: $\rho=\frac{\|\mathsf{A(u-u^{N_{x}+N_{y}-1})}\|_{\ell^{2}}}{\|\mathsf{A(u-u^{0})}\|_{\ell^{2}}}$
(b) RMS-PML: $\rho=\frac{\|\mathsf{A(u-u^{4})}\|_{\ell^{2}}}{\|\mathsf{A(u-u^{0})}\|_{\ell^{2}}}$
Figure 1: RAS-PML/RMS-PML on Checkerboard: The rate of reduction of residual against $k$. Left: Reduction rate per $N_{x}+N_{y}-1$ iterations for RAS-PML; Right: Reduction rate per exhaustive sweep for RMS-PML
N 2 4 8
@@
(c) RMS-PML: Left: strip $D$, Right checkerboard $D$.
Table 1: Iteration counts (relative residual tol = $10^{-6}$) with $f_{\rm PML}(x):=\frac{a}{3}x^{3}$, $a=30k$. The numbers without bracket are for the fixed point iterations and the numbers in brackets are for the preconditioned GMRES iterations. Note that $\frac{1}{80h}=4,7,10,13,16,19$ for $k=100,150,200,250,300,350$
Acknowledgement JG was supported by EPSRC grant EP/V001760/1. SG was supported by the National Natural Science Foundation of China (Grant number 12201535), the Guangdong Basic and Applied Basic Research Foundation (Grant number 2023A1515011651) and Shenzhen Stability Science Program 2022. DL was supported by INSMI (CNRS) through a PEPS JCJC grant 2023. EAS was supported by EPSRC grant EP/R005591/1. IGG was supported for several collaborative visits by CUHK Shenzhen.